OpenStudy (anonymous):

What are the approximate solutions of 2x² – x + 10 = 0? a) -2, 2.5 B) -1.97, 2.47 C) -2.5, 2 D) no solution

OpenStudy (mathstudent55):

Metjod 1, try each choice in the equation. a) Does -2 work? Plug in -2 for x in left side of equation and evaluate left side. 2(-2)^2 -(-2) + 10 = 8 + 2 + 10 = 20 -2 does not work Does 2.5 work? 2(2.5)^2 - 2.5 + 10 = 2*6.25 - 2.5 + 10 = 20 2.5 does not work

OpenStudy (mathstudent55):

Choice a does not work

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so for the next 2 parts i got 14.9028 and 20.5. so the answer would have to be D right?

OpenStudy (mathstudent55):

Now you can do the same method for choice B 2x^2 - x + 10 = 0 Let x = -1.97 (choice b) 2(-1.97)^2 - (-1.97) + 10 = 7.76 + 1.97 + 10 = 19.73 -1.97 does not work

OpenStudy (mathstudent55):

You still need to check c.

OpenStudy (anonymous):

c i got 20.5. is that right?

OpenStudy (mathstudent55):

Let's try -2.5 2(-2.5)^2 - (-2.5) + 10 = 12.5 + 2.5 + 10 = 25 -2.5 does not work, so c also does not work, so d is correct

OpenStudy (anonymous):

k. thanks!